Can you help with these two?Ask Your Teacher 6 1 points KatzPSEf1 2.P.047 My Notes + A uniformly accelerating rocket is found to have a velocity of 13.0+-12 points KatzPSEf1 2.P.061 My Notes +Ask Your T In the movie Star Wars: The Empire Strikes Back, after being told that Dar

Ask Your Teacher 6 1 points KatzPSEf1 2.P.047 My Notes + A uniformly accelerating rocket is found to have a velocity of 13.0 m/s when its height is 7.00 m above the ground, and 1.90 s later the rocket is at a height of 55.0 m. What is the magnitude of its acceleration? m/s2 Need Help? Read It Show My Work (optional)
+-12 points KatzPSEf1 2.P.061 My Notes +Ask Your T In the movie Star Wars: The Empire Strikes Back, after being told that Darth Vader is his father, Luke Skywalker falls from a ledge in Cloud City (not on the Earth, so the magnitude of the free-fall acceleration is not necessarily 9.81 m/s2). Suppose he falls distance of 54.0 m in 3.5 s. Assuming he starts from rest, answer the fol lowing questions. Use an upward-pointing y axis. (Express your answers in vector form.) (a) What is Luke's velocity 3.5 s after he starts to fall? m/s (b) What is the constant acceleration due to gravity, experienced by Luke, on Cloud City? m/s2 Need Help? Read It
Answer #1

u= 13-0 m}s. initial t = = 1. 90 sech 4 57) hm 48 m h= ut+at ax (90)2 48 = 13xI-90+ 48-24.7 = La x3-61 2. 23.3= 1.805 x a a 2